OpenStudy (anonymous):

45 mi/h = ___ ft/s a. 66 c. 26.4 b. 15.6 d. 19.2

hero (hero):

\[45 \dot\ \frac{\cancel{\text{mi}}}{\cancel{\text{hr}}} \dot\ \frac{5280}{1} \dot\ \frac{\text{ft}}{\cancel{\text{mi}}} \dot\ \frac{1}{3600} \dot\ \frac{\cancel{\text{hr}}}{\text{s}} = \]

OpenStudy (anonymous):

Do you have the answer in simplest form? please